Consider the perfectly competitive market for Diesel. The aggregate demand for gasoline is

The aggregate demand for gasoline is

Q_d=100-p

While the aggregate supply is Q_s= 3p

Work out the equilibrium price and quantity.  



1
Expert's answer
2021-08-24T14:28:04-0400

The equilibrium price and quantity is achieved at a point where

"Q_d=Q_s\\100-p=3p\\100=3p=p\\100=4p\\p=\frac{100}{4}\\p=25"


"Q_d=100-p\\Q_d=100-25\\Q_d=75"


"Q_s= 3p\\ Q_s= 3\times25\\ Q_s= 75"


Therefore equilibrium price=25 and equilibrium quantity is 75